how many liters of water must be added to 30L of a 20% chlorine solution to dilute it to a 15% chlorine solution
Answer by richwmiller(17219)

point of view of water
x+.80*30=.85(x+30)
x+24. = 0.85 (x+30)
x+24. = 0.85 x+25.5
0.15 x=1.5
x = 10
from the point of view of chlorine
.20*(30)=.15(x+30)
x = 10